A physical therapist wants to determine the difference in the proportion of men and women who participate in regular sustained physical activity. What sample size should be obtained if she wishes the estimate to be within five percentage points with 99% confidence, assuming that 


(a) she uses the estimates of 22.3% male and 18.9% female from a previous year? 

(b) she does not use any prior estimates?


 (a) n = _______ 
Answer #1

Margin of error, E = 0.05

Confidence Level, CL = 0.99

Significance level, α = 1 - CL = 0.01

Critical value, z = NORM.S.INV(0.01/2) =2.58

a) Proportion for sample 1, p̂₁ = 0.223  

Proportion for sample 2, p̂₂ = 0.189

Sample size, n = (z² * (p̂₁*(1-p̂₁) + p̂₂*(1-p̂₂)) / E²

= 869.459 = 869

n=869

b) Let p̂₁ = 0.50

p̂₂ = 0.50

Sample size, n = (z² * (p̂₁*(1-p̂₁) + p̂₂*(1-p̂₂)) / E²

= 1331.28= 1331

n=1331